# Understanding and Mitigating Errors of LLM-Generated RTL Code

### v1

@@ III-B Misinterpretation of Design Specifications
III-B1 Insufficient Understanding of Circuit Concepts (IUCC.)
@@
Inadequate Understanding of Timing Related Concepts (TRC.): In this scenario, the LLM’s lack of domain-specific knowledge in timing-related concepts leads to functional errors in the generated RTL code. These errors include mishandling synchronous/asynchronous logic, improper previous signal latching, and incorrect timing distinctions in assign vs. always block assignments. As shown in Fig. 3(a), the LLM incorrectly sets the sensitivity signals in the always block due to a lack of understanding of timing concepts in synchronous/asynchronous programming. Additionally, as illustrated in Fig. 3(b), because the LLM lacks knowledge about tracking the predecessor signal (prev_in) in edge detection, it incorrectly clears prev_in within the reset block. This prevents the module from correctly recording the previous state of in after reset, leading to functional errors.
@@
Limitations in Numerical and Vector Processing Capabilities (NVP.): Additionally, due to potential limitations in its understanding of numerical and vector processing, the LLM struggles to align high-level natural language descriptions with low-level RTL code, leading to errors. Typical examples of such errors include incorrect sign-bit handling logic, misconfigured case-condition vectors, and misunderstandings of least-significant-bit (LSB) concepts—all of which stem from deficiencies in numerical and vector comprehension. As shown in Fig 5, due to LLMs’ limited understanding of vector fuzzy matching (casez), when processing input vectors containing multiple ’1’s, the generated RTL code incorrectly selects the default branch, leading to functional errors. Additionally, the LLM-produced code contains numerical representation errors, such as invalid binary formats like 3’b2, where digits other than 0 and 1 are incorrectly included, further causing failures.
@@
We analyzed the distribution of sub-error categories under the ’Insufficient Understanding of Circuit Concepts’ scenario (Fig. 7), revealing that GPT-4 Turbo consistently demonstrated the lowest error rates across all subcategories, while GPT-3.5 Turbo showed the highest frequency of errors; Qwen-Coder-32B-Instruction’s performance fell between GPT-3.5 Turbo and GPT-4 Turbo. The analysis further reveals that both Qwen-Coder-32B and GPT-3.5 Turbo show significantly weaker comprehension of temporal concepts (particularly synchronous/asynchronous understanding) compared to GPT-4 Turbo, while all models demonstrate similar performance in special-design programming and state machine design for complex scenarios, indicating these knowledge gaps represent critical bottlenecks limiting LLMs’ programming accuracy.
Figure 8: Examples of ambigious design descriptions
III-B2 Ambiguous Design Descriptions (ADD.)
Unclear Overall Module Functionality (UOMF.): This type of error primarily manifests as internal contradictions in the overall functional description of the module or inconsistencies between the description and the Golden Design implementation. As shown in Fig. 8(a), the first part of the design description requires detecting any edge, while the end specifies capturing only rising edges (0 to 1). Due to this inherent contradiction, the LLM fails to correctly interpret the user’s intent, leading to design errors.
Ambiguous Module Input/Output Description (AIOD.): This type of error manifests as insufficient description of the basic functionality of each input/output signal. As shown in Fig. 8(b), the design description does not clearly specify the triggering mechanism of ena, leading the LLM to incorrectly assume it is edge-triggered rather than level-triggered, thereby introducing functional errors.
Missing Module Initialization Configuration (MMI.): Some design errors are caused by the description omitting module initialization information. As shown in Fig. 8(c), due to missing information in the description, the LLM-generated code lacks the initialization module compared to the Golden Design, resulting in functional errors.
Figure 9: Distribution of error subtypes under ADD.
We analyzed the distribution of subcategory errors under ambiguous design descriptions, as shown in Fig 11. The results indicate that, regardless of model capability, all models exhibit similar performance across different error categories. Comparatively, GPT-4 Turbo performs slightly better than the other two models, suggesting it generates marginally more robust code in scenarios with amibiguous descriptions.
Figure 10: Examples of multimodal data in hardware circuit design
III-B3 Misinterpretation of Multimodal Data (MMD.)
Figure 11: Distribution of error subtypes under MMD.
We analyzed the distribution of subcategory errors caused by misinterpretation of multimodal information. The results show that all LLMs performed poorly in comprehending multimodal information, with each model exhibiting over 25 design errors attributable to such misinterpretations.
Notably, GPT-4 Turbo demonstrated significantly better performance in parsing tabular data compared to GPT-3.5 Turbo and Qwen-Coder-32B-Instruction. However, for Karnaugh maps and state transfer diagrams, all models showed comparable performance. The identical error rates on KMAPs suggest that neither Qwen nor GPT series models have received specialized training on this data type.
Interestingly, in interpreting textual waveform descriptions (textual format), Qwen-Coder-32B-Instruction slightly outperformed GPT-4 Turbo, indicating that domain-specialized smaller models may possess capabilities absent in general large language models.
@@ III-B4 Missing Details of Long Descriptions (MDLD)
Although recent LLMs claim to support long-context understanding, we observe that they often focus primarily on the main instructions in the design descriptions, while overlooking finer details. While such omissions may not significantly impact general question-answering tasks, they pose substantial challenges in LLM-based code generation, where missing even minor details can result in functional errors. As illustrated in Fig. 12, the generated code fails to include the bit selection logic for out_difference, leading to incorrect behavior eventually.
@@ III-B5 Error Distribution Statistics Under Misinteroretation of Design Specifications
Our analysis reveals that for all evaluated LLMs, the primary bottleneck in generating correct RTL code stems from insufficient circuit knowledge leading to design description misinterpretations. This is particularly pronounced in Qwen-Coder-32B and GPT-3.5 Turbo, where over 40 erroneous designs originated from this limitation. In contrast, GPT-4 Turbo demonstrates significantly better comprehension of circuit programming concepts, suggesting that scaling model parameters and training data - particularly to enhance circuit knowledge - effectively reduces RTL generation errors.
A substantial portion of errors also arises from multimodal information processing failures. The comparable performance across models in this category indicates that neither Qwen nor GPT series models have received adequate cross-modal alignment training specific to hardware design domains.
Additionally, a smaller but consistent error category involves ambiguous descriptions and limited long-context understanding. Performance in this aspect remains similar across model scales, from small models to larger GPT variants. Enhancing LLMs’ robustness in generating RTL from vague specifications and improving their long-context comprehension in hardware design contexts remain critical research challenges.
@@ III-C Combination of Multiple Errors
As shown in Fig 14, the LLM initially sets the synchronously triggered reset as asynchronously triggered due to inadequate grasp of timing related concepts. Second, because of its limited understanding of numerical and vector concepts, when handling data transmitted in least-significant-first order, the LLM incorrectly configures the data recording direction as left-shift instead of the required right-shift. Additionally, due to deficiencies in RTL-specific programming knowledge, it mistakenly assigns a value to ’done’ (a wire type) inside an always block, leading to compilation errors. Finally, in complex state machine design scenarios, the LLM fails to account for error handling in the data transmission protocol. Upon reaching the STOP state, it only considers transitioning to DONE when the input is 1, neglecting the case where the input is 0—which requires waiting for an additional stop bit before returning to the START state. We attribute this error to the LLM’s limited knowledge of state machine design for complex scenarios. For such erroneous designs, we manually categorize them under each of these root causes simultaneously.
@@ IV-A Error Correction Mechanisms
Figure 15: The workflow of RAG-based knowledge error mitigation
Figure 16: An example of circuit design knowledge entry
RAG-based Knowledge Error Mitigation: We leverage RAG to supplement missing knowledge in both RTL programming and circuit design, and enhance the RTL code generation quality. To this end, we construct a domain‑specific knowledge base whose entries each consist of a keyword, a concise description, and an illustrative example. Circuit design knowledge is sourced from GitHub and StackOverflow and organized into four categories—arithmetic, memory, control, and miscellaneous modules—to enable efficient lookup (see Fig. 16).
@@
Multimodal Data Conversion: To mitigate generation errors arising from multimodal design inputs—such as Karnaugh maps, waveform diagrams, and truth tables, we have developed tools that convert these heterogeneous formats into a single, unified truth‑table meta‑representation as shown in Fig. 18. This meta‑representation can then be directly translated into Verilog via LLMs. Although the initially generated circuits may not be fully optimized, they can be automatically streamlined using logic synthesis tools under the specified design constraints. Importantly, when processing waveform inputs for sequential circuits, our conversion tool captures the clock triggering semantics to preserve correct timing behavior. Likewise, for state transition diagrams, the intermediate representation explicitly encodes both current and next states in addition to the input and output signals.
@@ IV-B Application of the Error Correction Mechanisms
We integrate our error‑correction mechanisms into a standard LLM‑based design workflow that comprises code generation, compilation, and verification as shown in Fig. 20. Before code generation, our RAG‑based mechanism retrieves relevant circuit designs and RTL programming knowledge from a domain‑specific knowledge base to enrich and disambiguate the user’s design descriptions. During the correction phase, compilation‑error messages similarly drive focused lookups to fill any remaining knowledge gaps. Concurrently, a rule‑guided refinement module applies design rule checks to the pre‑generation descriptions, automatically refining or repairing violations as needed. Whenever multimodal inputs (e.g., K‑maps, waveforms) are detected, a dedicated converter replaces them with our unified truth‑table meta‑representation, simplifying subsequent code generation. Finally, residual errors such as missing long‑context details are exposed via simulation. Our debugger then iteratively localizes faults and invokes LLM chain‑of‑thought reasoning to propose and apply fixes until all test cases pass. This integrated framework ensures robust and accurate RTL generation without manual intervention.
@@ V-A Experiment Setups
LLM Setups: We evaluate Verilog code generation and repair performance using GPT-based models. All experiments are conducted with a fixed temperature of 0.1. For the multi-round iterative RTL code repair method, we limit the maximum number of iterations to 10.
@@
Baselines: We compare against two categories of baselines: 1) Base models — single-pass RTL code generation approaches using both general LLMs (e.g., GPT-4 Turbo [31]), Code-Specific models (e.g. Qwen-Coder-32B-Instruction [32]) and RTL-specific models (e.g., ITERTL [33], CodeV [34]); 2) Agent-based systems — multi-agent frameworks designed to enhance LLM-based RTL generation, including LLM+RTLFixer [13], VeriAssistant [17], PromptV [25].
Methods Integration: In our experiment, we first refine all design descriptions using rule-based method to facilitate initial RTL generation or error correction. If the design still contains errors, we perform a single-round correction via multimodal data conversion if multimodal data is detected in the design description; otherwise, we apply a single-round correction using the RAG method. If errors persist, we then adopt the two-stage debugging mechanism for iterative repair.
@@ V-B End-to-End Code Generation Comparison
Table II presents a comparison between our method and existing baseline approaches. Our solution achieves 91.0% and 90.4% accuracy in pass@1 evaluations on VerilogEval-Human and VerilogEval-v2, respectively. Compared to strong vanilla models such as GPT-4 Turbo, our method demonstrates significant improvements of 32.7% and 29.5% on the two benchmarks. In addition, our approach outperforms various agent-based systems including LLM+RTLFixer, VeriAssistant, and PromptV. Meanwhile, our method achieves approximately a 30% improvement even on smaller-scale models like Qwen2.5-Coder-32B-Instruction and GPT-3.5 Turbo.
TABLE II: Main results
@@
Code-Specified LLM Qwen-Coder-32B-Instruction 46.8 46.1
@@
Ours Qwen-Coer-32B-Instruction 76.9 (+30.1) 78.1 (+32.0)
@@
GPT-4 Turbo 91.0 (+32.7) 90.4 (+29.5)
@@ V-C Error Correction Analysis
When utilizing foundational LLMs such as Qwen2.5-Coder-32B-Instruction, GPT-3.5 Turbo and GPT-4 Turbo to generate code on the VerilogEval-Human benchmark, we observe 243 erroneous RTL designs, respectively. Based on the integration strategy mentioned above, we use each model to self-correct their own erroneous designs. The number of successfully repaired designs for each method is summarized in Fig 21.
Impact of RAG-based Knowledge Error Mitigation (RAG.): Experimental results demonstrate that this method significantly reduces knowledge-related errors in RTL designs generated by GPT models. Specifically, it successfully corrects 22 erroneous designs for GPT-3.5 Turbo and 25 for GPT-4 Turbo. For the Qwen-Coder-32B-Instruction model, it successfully corrected 16 erroneous designs. These findings underscore that knowledge gaps remain a primary bottleneck for large language models in RTL code generation.
@@
Impact of Rule-based Description Refinement (RDR.): We observe that some design descriptions are vague, such as missing explanations of input signals. Additionally, we identify several issues stemming from illegal assignments and duplicate variable definitions. By clarifying descriptions and enforcing RTL programming conventions, this method enables GPT-3.5 Turbo and GPT-4 Turbo to correctly generate 9 and 6 additional designs, respectively. Using this approach, Qwen-Coder-32B-Instruction corrected 16 design errors. This reveals that compared to general-purpose LLMs, incorporating proper coding rules brings more significant improvements to models specifically pretrained on code.
Impact of Multimodal Data Conversion (MDC.): Our experiments reveal that current LLMs encounter challenges in semantically understanding multimodal hardware design data (e.g., Karnaugh maps). However, our tool-based conversion approach mitigates these issues, reducing errors by 13 instances in both GPT-3.5 Turbo and GPT-4 Turbo. For Qwen-Coder-32B-Instruction, this approach reduced 12 such errors.
@@
Impact of Two-Stage Debugging Mechanism (TDM.): We find that the debugging module performs better with GPT-4 Turbo than with GPT-3.5 Turbo and Qwen-Coder-32B-Instruction. This may be attributed to models like GPT-3.5 Turbo tendency to produce larger and more error-prone code blocks that require extensive revisions. Benefiting from enhanced reasoning and programming capabilities, GPT-4 Turbo exhibits greater precision in identifying design flaws and systematically correcting faulty RTL code.
TABLE III: Correction outcomes for different categories of errors
Error Category Qwen-Coder-32B-Instruction GPT-3.5 Turbo GPT-4 Turbo
Total Corrected Accuracy(%) Total Corrected Accuracy(%) Total Corrected Accuracy(%)
IKSP. 13 8 61.5 25 14 56.0 8 7 87.5
IUCC. 45 23 51.1 43 18 41.9 23 18 78.3
ADD. 10 3 30.0 11 3 27.3 8 4 50.0
MMD. 27 18 66.7 28 19 67.9 23 17 73.9
MDLD. 10 1 10.0 11 5 45.5 10 4 40.0
@@
Subsequently, we investigated the effectiveness of error correction across different error categories, based on 243 erroneous RTL designs generated by the three different LLMs above. The correction outcomes for each error category are summarized in Table III.
@@
It can be observed that for cases of Insufficient Knowledge of Specialized RTL Programming, the repair results are notably effective, with GPT-4 Turbo successfully fixing 87.5% of design failures caused by this issue. Regarding Insufficient Understanding of Circuit Concepts, all three models achieved repairs for approximately 20 faulty designs by incorporating additional external knowledge entries. When user descriptions are ambiguous, the framework can still rectify a subset of resulting errors, contributing to more robust code generation. For Misinterpretation of Multimodal Data, our method consistently repaired around 18 error-inducing designs across all three LLMs. In long-context scenarios, Qwen-Coder-32B-Instruction repaired one design, while GPT-3.5 Turbo and GPT-4 Turbo each fixed approximately five designs, indicating that general-purpose LLMs represented by the GPT series may demonstrate better long-context handling capabilities than specifically trained smaller models.

### v2

@@ III-B Misinterpretation of Design Specifications
III-B1 Ambiguous Design Descriptions (ADD.)
Figure 2: Examples of ambigious design descriptions
Unclear Overall Module Functionality (UOMF.): This type of error primarily manifests as internal contradictions in the overall functional description of the module or inconsistencies between the description and the Golden Design implementation. As shown in Fig. 2(a), the first part of the design description requires detecting any edge, while the end specifies capturing only rising edges (0 to 1). In the benchmark, the golden design implements ”detect any edge”, while the LLM implements ”0 to 1 transition occurs”. The ambiguity in the overall functional description within natural language leads to the LLM-generated design being evaluated as incorrect during evaluation.
Ambiguous Module Input/Output Description (AIOD.): This type of error manifests as insufficient description of the basic functionality of each input/output signal. As shown in Fig. 2(b), the design description does not clearly specify the triggering mechanism of ena. In the benchmark, the golden implementation is edge-triggered, while the LLM-generated design is level-triggered. This discrepancy leads to the LLM’s output being judged as a functional error during evaluation.
Missing Module Initialization Configuration (MMI.): Some design errors are caused by the description omitting module initialization information. As shown in Fig. 2(c), since the description did not specify initialization requirements, the LLM implemented the design strictly according to the given description. However, due to extra initialization in the golden design, the LLM-generated design was incorrectly judged as functionally incorrect during evaluation.
Figure 3: Distribution of error subtypes under ADD.
We analyzed the distribution of subcategory errors under ambiguous design descriptions, as shown in Fig 5. The results indicate that, regardless of model capability, all models exhibit similar performance across different error categories.
Figure 4: Examples of multimodal data in hardware circuit design
III-B2 Misinterpretation of Multimodal Data (MMD.)
Figure 5: Distribution of error subtypes under MMD.
We analyzed the distribution of subcategory errors caused by misinterpretation of multimodal information. The results show that all LLMs except Deepseek-v3.2-Speciale performed poorly in comprehending multimodal information, with each model exhibiting over 20 design errors attributable to such misinterpretations. Due to its excellent comprehension and reasoning capabilities, DeepSeek-V3.2-Speciale generated only two erroneous designs when provided with Waveform format inputs.
III-B3 Insufficient Understanding of Circuit Concepts (IUCC.)
@@
Inadequate Understanding of Timing Related Concepts (TRC.): In this scenario, the LLM’s lack of domain-specific knowledge in timing-related concepts leads to functional errors in the generated RTL code. These errors include mishandling synchronous/asynchronous logic, improper previous signal latching, and incorrect timing distinctions in assign vs. always block assignments. As illustrated in Fig. 6, because the LLM lacks knowledge about tracking the predecessor signal (prev_in) in edge detection, it incorrectly clears prev_in within the reset block. This prevents the module from correctly recording the previous state of in after reset, leading to functional errors.
@@
Limitations in Numerical and Vector Processing Capabilities (NVP.): Additionally, due to potential limitations in its understanding of numerical and vector processing, the LLM struggles to align high-level natural language descriptions with low-level RTL code, leading to errors. Typical examples of such errors include incorrect sign-bit handling logic, misconfigured case-condition vectors, and misunderstandings of least-significant-bit (LSB) concepts—all of which stem from deficiencies in numerical and vector comprehension.
As shown in Fig 8, the design description requires outputting the position of the lowest ’1’ bit in the input vector (e.g., for input 8’b10010000, output 4). The correct design uses casez with wildcard matching to locate this bit. However, due to a lack of knowledge in processing vectors and numbers, the LLM incorrectly uses exact case matching. This fails for inputs with multiple ’1’s (e.g., 8’b10010000 is misrouted to the default branch, outputting 0), resulting in a functional error.
@@
We have statistically analyzed the distribution of various error subtypes under the scenario of ”Insufficient Understanding of Circuit Concept”, as shown in Figure 6. It can be observed that, with the increase in model scale and capabilities enhanced by reasoning, DeepSeek-V3.2-Speciale makes significantly fewer knowledge-related errors than the other four models. Meanwhile, Qwen2.5-Coder-32B-Instruct and GPT-3.5 Turbo exhibit notably more errors related to timing and vector/numerical processing compared to the other models. Increasing the model scale and incorporating domain-specific training data can effectively reduce these types of errors. However, all models show a relatively high number of errors in state machine design within complex circuits, indicating that this may be a key bottleneck limiting LLMs’ ability to correctly generate RTL designs.
@@ III-B4 Missing Details of Long Descriptions (MDLD)
Although recent LLMs claim to support long-context understanding, we observe that they often focus primarily on the main instructions in the design descriptions, while overlooking finer details. While such omissions may not significantly impact general question-answering tasks, they pose substantial challenges in LLM-based code generation, where missing even minor details can result in functional errors. As illustrated in Fig. 11, in the context of long input sequences, the design description requires that the i-th bit of out_different outputs the XOR value between the i-th bit of the in vector and its left neighbor. However, the LLM’s implementation overlooks this detail and incorrectly computes the XOR with the right neighbor instead, resulting in a functional error.
@@ III-B5 Error Distribution Statistics Under Misinteroretation of Design Specifications
Our analysis reveals that for all evaluated LLMs, the primary bottleneck in generating correct RTL code stems from insufficient circuit knowledge leading to design description misinterpretations. This is particularly pronounced in Qwen2.5-Coder-32B-Instruct and GPT-3.5 Turbo, where over 40 erroneous designs originated from this limitation. GPT-4 Turbo and CodeV-R1-RL-Qwen-7B also have more than 20 designs affected by such flaws, while Deepseek-v3.2 Speciale, due to its string programming capabilities, has only 6 related errors.
A substantial portion of errors also arises from multimodal information processing failures. The comparable performance across models in this category indicates that neither Qwen, CodeV-R1 nor GPT series models have received adequate cross-modal alignment training specific to hardware design domains. Thanks to its powerful comprehension and reasoning capabilities, Deepseek-v3.2-Speciale far surpasses other models in addressing such issues.
Additionally, a smaller but consistent error category involves ambiguous descriptions and limited long-context understanding. Performance in this aspect remains similar across model scales, from basic general-purpose models to powerful reasoning models. Enhancing LLMs’ robustness in generating RTL from vague specifications and improving their long-context comprehension in hardware design contexts remain critical research challenges.
@@ III-C Combination of Multiple Errors
  • Error 1: Inadequate Understanding of Timing Related Concepts: The user requested the implementation of a synchronous triggering circuit, but the LLM generated an asynchronous triggering circuit.
  • Error 2: Limitations in Numerical and Vector Processing Capabilities: The user specified Least Significant Bit first data input. To record data correctly, the data register should first store the incoming bit at the highest bit position and then shift right. However, the LLM implemented the shift in the opposite direction.
  • Error 3: Illegal Assignment to Wire in Always Block: The LLM incorrectly assigned a value to a wire-type variable within an always block, which violates RTL programming rules. This failure is attributed to the LLM’s deficient grasp of coding conventions.
  • Error 4: Deficiencies in State Machine Design for Complex Scenarios: Upon reaching the data transmission completion state, LLM’s designed state machine lacks the state transition logic for error handling.
@@ IV-A Error Correction Mechanisms
Multimodal Data Conversion: To mitigate generation errors arising from multimodal design inputs—such as Karnaugh maps, waveform diagrams, and truth tables, we have developed tools that convert these heterogeneous formats into a single, unified truth‑table meta‑representation as shown in Fig. 15.
Specifically, we find that although elements such as Karnaugh maps and waveforms are difficult for LLM to understand, their inputs adhere to specific norms and standards. For instance, Karnaugh maps inputs include variable definitions in rows/columns and corresponding 0/1 arrangements, while waveforms follow a tabular format where the first column of each row contains clock information. Therefore, we first extract the multimodal information from descriptions using regular expression matching. Subsequently, we parse the extracted information based on the applicable input norms, determine the input conditions corresponding to each output item, and map them to the meta-representation of a truth table. LLMs can then translate this meta-representation into the corresponding RTL design.
Importantly, when processing waveform inputs for sequential circuits, our conversion tool captures the clock triggering semantics to preserve correct timing behavior. Likewise, for state transition diagrams, the intermediate representation explicitly encodes both current and next states in addition to the input and output signals.
Figure 16: The workflow of RAG-based knowledge error mitigation
RAG-based Knowledge Error Mitigation: We leverage RAG to supplement missing knowledge in both RTL programming and circuit design, and enhance the RTL code generation quality. To this end, we construct a domain‑specific knowledge base whose entries each consist of a keyword, a concise description, and an illustrative example.
The knowledge base comprises two primary components. Circuit design knowledge entries are constructed through a three-stage pipeline: high-quality code is first collected from GitHub using keywords such as “RTL tutorial” and “arithmetic circuit”; the collected repositories are then filtered by excluding those with fewer than 100 stars and heuristically removing non-functional text like licenses to reduce noise; finally, the code is annotated via few-shot learning with GPT-4 Turbo to generate natural language functional descriptions, from which relevant keywords are extracted. We have collected approximately 500 entries covering various types of circuits such as arithmetic and memory modules. Separately, RTL programming knowledge entries are manually curated to address common syntax errors, with each entry providing error message, repair suggestions and concrete correction examples. We have manually designed 15 such entries. An illustrative entry is shown in Fig 17.
Figure 17: An example of circuit design knowledge entry
@@ IV-B Application of the Error Correction Mechanisms
We integrate our error‑correction mechanisms into a standard LLM‑based design workflow that comprises code generation, compilation, and verification as shown in Fig. 19.
Given a description, the rule-based description refinement module first examines and clarifies the description while incorporating necessary RTL programming specifications to achieve initial RTL generation. If the generated RTL still contains errors, multimodal data from the description is extracted and converted via multimodal data conversion. A dedicated converter replaces such data with a unified truth-table meta-representation, simplifying subsequent code generation. If issues persist or no multimodal input is detected, our RAG-based mechanism retrieves relevant circuit designs and RTL programming knowledge from a domain-specific knowledge base to enrich the user’s design descriptions, reducing design faults due to knowledge gaps. Finally, residual errors such as missing long-context details are exposed through simulation. Our debugger then iteratively localizes faults and invokes LLM chain-of-thought reasoning to propose and apply fixes until all test cases pass. This integrated framework ensures robust and accurate RTL generation without manual intervention.
@@ V-A Experiment Setups
LLM Setups: We evaluate Verilog code generation and repair performance using GPT-based models. All experiments are conducted with a fixed temperature of 0.1, n = 1. For the multi-round iterative RTL code repair method, we limit the maximum number of iterations to 10.
@@
Baselines: We compare against two categories of baselines: 1) Base models — single-pass RTL code generation approaches using both General LLMs (e.g., Deepseek-v3.2-Speciale), Code-Specific models (e.g. Qwen2.5-Coder-32B-Instruct[7]) and RTL-specific models (e.g., CodeV-R1-RL-Qwen-7B); 2) Agent-based systems — multi-agent frameworks designed to enhance LLM-based RTL generation, including MAGE [29], LLM+RTLFixer [24], PromptV [13].
Methods Integration: In our experiment, we first refine all design descriptions using rule-based method to facilitate initial RTL generation or error correction. If the design still contains errors, we perform a single-round correction via multimodal data conversion if multimodal data is detected in the design description; then we apply a single-round correction using the RAG method. If errors persist, we then adopt the two-stage debugging mechanism for iterative repair.
@@ V-B End-to-End Code Generation Comparison
TABLE III: Main results
@@
GPT-3.5 Turbo 42.9 41.0
Code-Specified LLM Qwen2.5-Coder-32B-Instruct 46.8 46.1
@@
MAGE[29] Claude 3.5 Sonnet 89.1 93.6
Ours Deepseek-v3.2-Speciale 98.1 (+9.0) 97.4 (+9.6)
GPT-4 Turbo 91.0 (+32.7) 90.4 (+29.5)
@@
Qwen2.5-Coder-32B-Instruct 76.9 (+30.1) 78.1 (+32.0)
CodeV-R1-RL-Qwen-7B 89.7 (+23.0) 91.7 (+23.9)
Table III presents a comparison between our method and existing baseline approaches. Our solution achieves 98.1% and 97.4% accuracy in pass@1 evaluations on VerilogEval-Human and VerilogEval-v2, respectively. As can be seen, our method delivers performance improvements ranging from approximately 9% to 30%, whether applied to the powerful general-purpose model DeepSeek-v3.2-Speciale or to smaller models additionally trained on Code and RTL. In addition, our approach outperforms various agent-based systems including MAGE,LLM+RTLFixer, and PromptV. Meanwhile, by integrating our code generation and repair framework, CodeV-R1-RL-Qwen-7B achieves performance comparable to the current SOTA agent framework, MAGE. Meanwhile, based on DeepSeek-v3.2-Speciale, our framework demonstrates performance improvements over MAGE of 9.0% on VerilogEval-Human and 4.1% on VerilogEval-v2, respectively.
@@ V-C Error Correction Analysis
When utilizing foundational LLMs to generate code on the VerilogEval-Human benchmark, we use each model to self-correct their own erroneous designs, as shown in Table I. The number of successfully repaired designs for each method is summarized in Fig 20.
@@
Impact of Rule-based Description Refinement (RDR.): We observe that some design descriptions are vague, such as missing explanations of input signals. Additionally, we identify several issues stemming from illegal assignments. It can be seen that all models based on this module can fix more than 6 faulty designs, and in particular, Qwen2.5-Coder-32B-Instruct can fix 16 designs.
Impact of Multimodal Data Conversion (MDC.): Our experiments reveal that current LLMs encounter challenges in semantically understanding multimodal hardware design data. It can be seen that all models based on this module—except for Deepseek-v3.2-Speciale model—can correctly fix 9 or more designs. Moreover, integrating this module into Deepseek-v3.2-Speciale can also fully fix all errors caused by misunderstandings of Waveform.
Impact of RAG-based Knowledge Error Mitigation (RAG.): Experimental results demonstrate that this method significantly reduces knowledge-related errors in RTL designs generated by GPT models. Meanwhile, this module can also reduce 15 faults caused by deficiencies in circuit knowledge mastery in Qwen2.5-Coder-32B-Instruct and CodeV-R1. On Deepseek-v3.2-Speciale, this module can also fix approximately 4 faulty designs.
@@
Impact of Two-Stage Debugging Mechanism (TDM.): We find that the debugging module performs better with GPT-4 Turbo and CodeV-R1 than with GPT-3.5 Turbo and Qwen2.5-Coder-32B-Instruct. This may be attributed to models like GPT-3.5 Turbo tendency to produce larger and more error-prone code blocks that require extensive revisions.
TABLE IV: Correction outcomes for different categories of errors
Model IKSP. ADD. MMD. IUCC. MDLD.
Total Pass Acc Total Pass Acc Total Pass Acc Total Pass Acc Total Pass Acc
Deepseek-v3.2-Speciale 2 2 100.0 5 4 80.0 2 2 100.0 6 5 83.3 3 2 66.7
GPT-4 Turbo 8 7 87.5 8 4 50.0 23 17 73.9 23 18 78.3 10 4 40.0
GPT-3.5 Turbo 25 14 56.0 11 3 27.3 28 19 67.9 43 18 41.9 11 5 45.5
Qwen2.5-Coder-32B-Instruct 13 8 61.5 10 3 30.0 27 18 66.7 45 23 51.1 10 1 10.0
CodeV-R1-RL-Qwen-7B 3 2 66.7 5 3 60.0 23 13 56.5 22 16 72.7 11 4 36.4
*Acc: Accuracy (%), calculated as Pass/Total × 100.
@@
Subsequently, we investigated the effectiveness of error correction across different error categories, based on 243 erroneous RTL designs generated by the three different LLMs above. The correction outcomes for each error category are summarized in Table IV.
@@
It can be observed that for cases of Insufficient Knowledge of Specialized RTL Programming, the repair results are notably effective, On all models, this method can fix over 60% of such incorrectly designed cases. When user descriptions are ambiguous, the framework can still rectify a subset of resulting errors, contributing to more robust code generation. For Misinterpretation of Multimodal Data, our framework can fix more than 13 incorrectly designed cases on all models except Deepseek-v3.2-Speciale. Similarly, for Insufficient Understanding of Circuit Concepts, our framework can fix over 16 incorrectly designed cases caused by this issue. Furthermore, in complex long-context scenarios, it can be observed that when our framework is based on general-purpose LLMs, its repair performance is better compared to code-specific LLMs—it can fix over 40% of the designs with faults caused by this scenario.